# Uniformly-$S$-pseudo-projective modules

### v2

@@ 2. $u$-$S$-pseudo-projective modules
Let $R=\mathbb{Z}$ and $S=\{1,2,3,\cdots\}$. Let $\mathbb{P}$ be the set of all prime numbers. Then $\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}$ is a $u$-$S$-quasi-projective module that is not $u$-$S$-projective by [2, Example 3.3]. By Corollary 2.10, $\Big(\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}\Big)\oplus\Big(\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}\Big)$ is a $u$-$S$-pseudo-projective module. However, $\Big(\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}\Big)\oplus\Big(\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}\Big)$ is not $u$-$S$-projective by [2, Corollary 2.8 (1)].
@@
Let $S$ be a multiplicative subset of a ring $R$. The following statements hold:
  1. (1)
    If $A\oplus B$ is $u$-$S$-pseudo-projective, then any $u$-$S$-epimorphism $f:A\to B$ $u$-$S$-splits.
  2. (2)
    If $M$ is an $R$-module and $f:P\to M$ is an epimorphism with $P$ projective, then $M$ is $u$-$S$-projective if and only if $P\oplus M$ is $u$-$S$-pseudo-projective.
@@
  1. (1)
    Let $f:A\to B$ be any $u$-$S$-epimorphism. Since $A\oplus B$ is $u$-$S$-pseudo-projective, then by Proposition 2.9, $B$ is $u$-$S$-projective relative to $A$ and hence by [2, Lemma 2.15], the $u$-$S$-exact sequence $0\to\text{Ker}(f)\to A\xrightarrow{f}B\to 0$ $u$-$S$-splits. Thus $f$ $u$-$S$-splits.
  2. (2)
    Suppose that $M$ is $u$-$S$-projective. Since $P$ is projective, then by [9, Corollary 2.11], $P$ is $u$-$S$-projective. So by [9, Proposition 2.14 (1)], $P\oplus M$ is $u$-$S$-projective and hence by Remark 2.2 (2), $P\oplus M$ is $u$-$S$-pseudo-projective. Conversely, suppose that $P\oplus M$ is $u$-$S$-pseudo-projective. Then by part (1) and since $f:P\to M$ is a $u$-$S$-epimorphism, we have $f$ $u$-$S$-splits. So the exact sequence $0\to\text{Ker}(f)\to P\xrightarrow{f}M\to 0$ $u$-$S$-splits. Hence by [5, Lemma 2.8], $P$ is $u$-$S$-isomorphic to $\text{Ker}(f)\oplus M$. Since $P$ is $u$-$S$-projective, then by [9, Proposition 2.14 (3)] and [2, Corollary 2.8 (1)], $M$ is $u$-$S$-projective.
@@
Let $M$ be $u$-$S$-injective. Then $M\oplus E(M)$ is $u$-$S$-injective by [6, Proposition 4.7]. So by [3, Remark 2.4 (2)], $M\oplus E(M)$ is $u$-$S$-pseudo-injective. Conversely, suppose that $M\oplus E(M)$ is $u$-$S$-pseudo-injective. Then by [3, Theorem 2.6], the monomorphism $i_{M}:M\to E(M)$ $u$-$S$-splits. So the exact sequence $0\to M\to E(M)\to\frac{E(M)}{M}\to 0$ $u$-$S$-splits. Hence by [5, Lemma 2.8], $E(M)$ is $u$-$S$-isomorphic to $M\oplus\frac{E(M)}{M}$. But $E(M)$ is $u$-$S$-injective, so by [6, Proposition 4.7 (3)] and [2, Corollary 2.8 (2)], $M$ is $u$-$S$-injective. ∎
@@
    Every $R$-module is $u$-$S$-pseudo-projective.
@@
    Every direct sum of an $R$-module and a free $R$-module is $u$-$S$-pseudo-projective ($u$-$S$-pseudo-injective).
@@
$(2)\Rightarrow(1)$. Let $M$ be any $R$-module and let $f:P\to M$ be an epimorphism with $P$ projective. Then by (2), $P\oplus M$ is $u$-$S$-pseudo-projective. So by Proposition 2.13 (2), $M$ is $u$-$S$-projective. Hence every $R$-module is $u$-$S$-projective. Therefore, $R$ is $u$-$S$-semisimple by [9, Theorem 3.5].
@@
$(3)\Rightarrow(1)$. Let $F$ be any free $R$-module. Then by (3), for any $R$-module $M$, $M\oplus F$ is $u$-$S$-pseudo-projective ($u$-$S$-pseudo-injective). By Proposition 2.9 ([3, Proposition 2.13]), $M$ is $u$-$S$-projective ($u$-$S$-injective) relative to $F$ for any $R$-module $M$. Thus every $R$-module is $u$-$S$-projective ($u$-$S$-injective) relative to $F$, By [2, Theorem 2.16], $F$ is $u$-$S$-semisimple. Hence every free $R$-module is $u$-$S$-semisimple. Therefore, $R$ is $u$-$S$-semisimple.
@@
    Every $u$-$S$-quasi-projective module is $u$-$S$-projective.
@@
    Every direct sum of two $u$-$S$-quasi-projective modules is $u$-$S$-projective.
@@
$(1)\Rightarrow(2)$. Let $M$ and $N$ be two $u$-$S$-quasi-projective modules. Then by (1), $M$ and $N$ are $u$-$S$-projective modules. So $M\oplus N$ is $u$-$S$-projective by [9, Proposition 2.14 (1)].
@@
commutes for some $s\in S$. So $sp_{2}=fp_{1}e$. Since $p_{2}i_{2}=1_{M}$, then $s1_{M}=sp_{2}i_{2}=fp_{1}ei_{2}$. Let $q:=p_{1}ei_{2}:M\to P$. Then $s1_{M}=fq$. So by [9, Lemma 2.4], the exact sequence $0\to\text{Ker}(f)\to P\xrightarrow{f}M\to 0$ $u$-$S$-splits. By [5, Lemma 2.8], $P$ is $u$-$S$-isomorphic to $\text{Ker}(f)\oplus M$. Since $P$ is $u$-$S$-projective, $\text{Ker}(f)\oplus M$ is $u$-$S$-projective by [9, Proposition 2.14 (3)]. Hence $M$ is $u$-$S$-projective by [2, Corollary 2.8 (1)]. Thus (1) holds. ∎
@@
$(3)\Rightarrow(1)$. Let $M$ be any $u$-$S$-pseudo-projective module. Let $f:P\to M$ be an epimorphism with $P$ projective. Then by (3), $P\oplus M$ is $u$-$S$-pseudo-projective. Thus $M$ is $u$-$S$-projective by Proposition 2.13 (2).
Let $\mathfrak{p}$ be a prime ideal of a ring $R$. Then $S=R\setminus\mathfrak{p}$ is a multiplicative subset of $R$. We say that an $R$-module $M$ is $u$-$\mathfrak{p}$-pseudo-projective if $M$ is $u$-$S$-pseudo-projective.
Proposition 2.18.
Let $S$ be a multiplicative subset of a ring $R$ and $P$ an $R$-module. Suppose that $P$ is $u$-$\mathfrak{m}$-pseudo-projective for every $\mathfrak{m}\in\text{Max}(R)$. Then $P$ is pseudo-projective.
Proof.
Let $K\leq P$ and $f:P\to\frac{P}{K}$ be any epimorphism. Then for every $\mathfrak{m}\in\text{Max}(R)$, there is $s_{\mathfrak{m}}\in S$ and $e_{\mathfrak{m}}\in\text{End}_{R}(P)$ such that $s_{\mathfrak{m}}f=\eta_{K}e_{\mathfrak{m}}$. Since $R=\langle\{s_{\mathfrak{m}}\mid\mathfrak{m}\in\text{Max}(R)\}\rangle$, so there are finite sets $\{r_{\mathfrak{m}_{i}}\}_{i=1}^{n}\subseteq R$ and $\{s_{\mathfrak{m}_{i}}\}_{i=1}^{n}\subseteq S$ such that $1=\sum\limits_{i=1}^{n}r_{\mathfrak{m}_{i}}s_{\mathfrak{m}_{i}}$. Let $e=\sum\limits_{i=1}^{n}r_{\mathfrak{m}_{i}}e_{\mathfrak{m_{i}}}$. Then $e\in\text{End}_{R}(P)$ and $f=\sum\limits_{i=1}^{n}r_{\mathfrak{m}_{i}}s_{\mathfrak{m}_{i}}f=\sum\limits_{i=1}^{n}r_{\mathfrak{m}_{i}}\eta_{K}e_{\mathfrak{m_{i}}}=\eta_{K}\sum\limits_{i=1}^{n}r_{\mathfrak{m}_{i}}e_{\mathfrak{m_{i}}}=\eta_{K}e$. Therefore, $P$ is pseudo-projective. ∎
Remark 2.19.
Let $S$ be a multiplicative subset of a ring $R$ and $M$ an $R$-module. If $M\oplus M$ is pseudo-projective, then $M\oplus M$ is $u$-$S$-pseudo-projective.
Proof.
Suppose that $M\oplus M$ is pseudo-projective. Then $M$ is quasi-projective by [7, Lemma 8]. Hence $M$ is $u$-$S$-quasi-projective by [2, Remark 3.2 (2)]. Thus $M\oplus M$ is $u$-$S$-pseudo-projective by Corollary 2.10. ∎
Proposition 2.20.
Let $S$ be a multiplicative subset of a ring $R$ and $P$ an $R$-module. Then the following statements are equivalent:
  1. (1)
    $P\oplus P$ is pseudo-projective.
  2. (2)
    $P\oplus P$ is $u$-$\mathfrak{p}$-pseudo-projective for every $\mathfrak{p}\in\text{Spec}(R)$.
  3. (3)
    $P\oplus P$ is $u$-$\mathfrak{m}$-pseudo-projective for every $\mathfrak{m}\in\text{Max}(R)$.
Proof.
$(1)\Rightarrow(2)$. Follows from Remark 2.19.
$(2)\Rightarrow(3)$. Clear.
$(3)\Rightarrow(1)$. Follows from Proposition 2.18. ∎
Example 2.21.
Since $\mathbb{Z}$ is a quasi-projective $\mathbb{Z}$-module, then $\mathbb{Z}\oplus\mathbb{Z}$ is quasi-projective and so $\mathbb{Z}\oplus\mathbb{Z}$ is pseudo-projective. By Proposition 2.20, $\mathbb{Z}\oplus\mathbb{Z}$ is $u$-$p\mathbb{Z}$-pseudo-projective for every prime number $p$.
Acknowledgments: The authors would like to thank the referees for their valuable comments.

### v3

@@ 2. $u$-$S$-pseudo-projective modules
Let $R=\mathbb{Z}$ and $S=\{1,2,3,\cdots\}$. Let $\mathbb{P}$ be the set of all prime numbers. Then $M:=\bigoplus\limits_{p\in\mathbb{P}}\mathbb{Z}_{p}$ is a $u$-$S$-quasi-projective module that is not $u$-$S$-projective by [2, Example 3.3]. By Corollary 2.10, $M\oplus M$ is a $u$-$S$-pseudo-projective module. However, $M\oplus M$ is not $u$-$S$-projective by [2, Corollary 2.8 (1)].
@@
Let $S$ be a multiplicative subset of a ring $R$, $M$ an $R$-module, and $f:P\to M$ an epimorphism with $P$ projective. Then $M$ is $u$-$S$-projective if and only if $P\oplus M$ is $u$-$S$-pseudo-projective.
@@
Suppose that $M$ is $u$-$S$-projective. Since $P$ is projective, then by [10, Corollary 2.11], $P$ is $u$-$S$-projective. So by [10, Proposition 2.14 (1)], $P\oplus M$ is $u$-$S$-projective and hence by Remark 2.2 (2), $P\oplus M$ is $u$-$S$-pseudo-projective. Conversely, suppose that $P\oplus M$ is $u$-$S$-pseudo-projective. Then by Proposition 2.9 (2) and since $f:P\to M$ is a $u$-$S$-epimorphism, we have $f$ $u$-$S$-splits. So the exact sequence $0\to\text{Ker}(f)\to P\xrightarrow{f}M\to 0$ $u$-$S$-splits. Hence by [6, Lemma 2.8], $P$ is $u$-$S$-isomorphic to $\text{Ker}(f)\oplus M$. Since $P$ is $u$-$S$-projective, then by [10, Proposition 2.14 (3)] and [2, Corollary 2.8 (1)], $M$ is $u$-$S$-projective. ∎
Corollary 2.14.
Let $R$ be a ring, $M$ an $R$-module, and $f:P\to M$ an epimorphism with $P$ projective. Then $M$ is projective if and only if $P\oplus M$ is pseudo-projective.
@@
Let $M$ be $u$-$S$-injective. Then $M\oplus E(M)$ is $u$-$S$-injective by [7, Proposition 4.7]. So by [3, Remark 2.4 (2)], $M\oplus E(M)$ is $u$-$S$-pseudo-injective. Conversely, suppose that $M\oplus E(M)$ is $u$-$S$-pseudo-injective. Then by [3, Theorem 2.6], the monomorphism $i_{M}:M\to E(M)$ $u$-$S$-splits. So the exact sequence $0\to M\to E(M)\to\frac{E(M)}{M}\to 0$ $u$-$S$-splits. Hence by [6, Lemma 2.8], $E(M)$ is $u$-$S$-isomorphic to $M\oplus\frac{E(M)}{M}$. But $E(M)$ is $u$-$S$-injective, so by [7, Proposition 4.7 (3)] and [2, Corollary 2.8 (2)], $M$ is $u$-$S$-injective. ∎
Corollary 2.16.
Let $R$ be a ring and $M$ an $R$-module. Then $M$ is injective if and only if $M\oplus E(M)$ is pseudo-injective.
Recall that an $R$-module $F$ is called $u$-$S$-flat if the induced sequence $0\to A\otimes_{R}F\to B\otimes_{R}F\to C\otimes_{R}F\to 0$ is $u$-$S$-exact for any $u$-$S$-exact sequence $0\to A\to B\to C\to 0$ [11], and that a ring $R$ is called a strongly $S$-perfect ring if every $u$-$S$-flat is projective [5].
Theorem 2.17.
  1. (1)
    $R$ is strongly $S$-perfect;
  2. (2)
    every $u$-$S$-flat is quasi-projective;
  3. (3)
    every $u$-$S$-flat is pseudo-projective.
Proof.
$(1)\Rightarrow(2)\Rightarrow(3)$. Clear.
$(3)\Rightarrow(1)$. Let $F$ be any $u$-$S$-flat $R$-module and $f:P\to F$ be an epimorphism with $P$ projective. Since $P$ is projective, it is $u$-$S$-flat by [10, Corollary 2.11 and Proposition 2.13]. So by [11, Proposition 3.4 (2)], $P\oplus F$ is $u$-$S$-flat. By (3), $P\oplus F$ is pseudo-projective, and so by Corollary 2.14, $F$ is projective. Hence, every $u$-$S$-flat $R$-module is projective. Therefore, $R$ is a strongly $S$-perfect ring. ∎
Theorem 2.18.
  1. (1)
    every finitely generated $u$-$S$-flat is projective;
  2. (2)
    every finitely generated $u$-$S$-flat is quasi-projective;
  3. (3)
    every finitely generated $u$-$S$-flat is pseudo-projective.
Proof.
Similar to the proof of Theorem 2.17. ∎
@@
    every $R$-module is $u$-$S$-pseudo-projective;
@@
    every direct sum of an $R$-module and a free $R$-module is $u$-$S$-pseudo-projective ($u$-$S$-pseudo-injective);
@@
$(2)\Rightarrow(1)$. Let $M$ be any $R$-module and let $f:P\to M$ be an epimorphism with $P$ projective. Then by (2), $P\oplus M$ is $u$-$S$-pseudo-projective. So by Proposition 2.13, $M$ is $u$-$S$-projective. Hence every $R$-module is $u$-$S$-projective. Therefore, $R$ is $u$-$S$-semisimple by [10, Theorem 3.5].
@@
$(3)\Rightarrow(1)$. Let $F$ be any free $R$-module. Then by (3), for any $R$-module $M$, $M\oplus F$ is $u$-$S$-pseudo-projective ($u$-$S$-pseudo-injective). By Proposition 2.9 (1) ([3, Proposition 2.13]), $M$ is $u$-$S$-projective ($u$-$S$-injective) relative to $F$ for any $R$-module $M$. Thus every $R$-module is $u$-$S$-projective ($u$-$S$-injective) relative to $F$, By [2, Theorem 2.16], $F$ is $u$-$S$-semisimple. Hence every free $R$-module is $u$-$S$-semisimple. Therefore, $R$ is $u$-$S$-semisimple.
@@
    Every $u$-$S$-quasi-projective module is $u$-$S$-projective;
@@
    Every direct sum of two $u$-$S$-quasi-projective modules is $u$-$S$-projective;
@@
$(1)\Rightarrow(2)$. Let $M$ and $N$ be two $u$-$S$-quasi-projective modules. Then by (1), $M$ and $N$ are $u$-$S$-projective modules. So $M\oplus N$ is $u$-$S$-projective by [10, Proposition 2.14 (1)].
@@
commutes for some $s\in S$. So $sp_{2}=fp_{1}e$. Since $p_{2}i_{2}=1_{M}$, then $s1_{M}=sp_{2}i_{2}=fp_{1}ei_{2}$. Let $q:=p_{1}ei_{2}:M\to P$. Then $s1_{M}=fq$. So by [10, Lemma 2.4], the exact sequence $0\to\text{Ker}(f)\to P\xrightarrow{f}M\to 0$ $u$-$S$-splits. By [6, Lemma 2.8], $P$ is $u$-$S$-isomorphic to $\text{Ker}(f)\oplus M$. Since $P$ is $u$-$S$-projective, $\text{Ker}(f)\oplus M$ is $u$-$S$-projective by [10, Proposition 2.14 (3)]. Hence $M$ is $u$-$S$-projective by [2, Corollary 2.8 (1)]. Thus (1) holds. ∎
@@
$(3)\Rightarrow(1)$. Let $M$ be any $u$-$S$-pseudo-projective module. Let $f:P\to M$ be an epimorphism with $P$ projective. Then by (3), $P\oplus M$ is $u$-$S$-pseudo-projective. Thus $M$ is $u$-$S$-projective by Proposition 2.13. ∎